# Rethinking Programmed I/O for Fast Devices, Cheap Cores, and Coherent Interconnects

### v2

@@ 5. PIO over a coherent interconnect
    The CPU cache receives line A’s new data and the CPU unblocks, immediately loading the first word of the invocation response into a register.
As described, the end state is that A is now Invalid and B is Shared in the CPU cache. However, in some deployed MESI-like protocols (including the Enzian interconnect we use), the device is free to return the data stipulating that it is not Shared but now Exclusive to the CPU. This final feature leaves the cache in the state it started in, with the roles of A and B reversed, ready for the next transaction.
The whole exchange takes 2 round trips on the coherent interconnect, and require no spinning on the CPU.
The other 3 protocols work similarly. For example, 5(b) shows the analogous case where the device initiates the request to the CPU, but still using lines homed on the device. This is used for the NIC experiment in Section 6.3.
The remaining two cases, where the cache lines are homed on the CPU, are more complex to deal with, since there is a possibility that the CPU cache will evict one of the cache lines without warning. The device will not observe this, and so might incur an access to main memory on the CPU node. In this paper we therefore focus on the case where the cache lines are homed on the device.
Subtleties:
A number of issues must be handled to make this practical, starting with timeouts: we block a core’s request for a cache line until a device operation has completed. A timeout in the CPU’s cache is likely to cause a memory fault exception, a fatal machine check, or the processor simply locking up (on Enzian, it is a machine check).
@@
The second issue concerns deadlock due to the device stalling a request until it can issue, and get a reply back for, another request to a different cache line. Existing standards for interoperable cache coherent interconnects are silent on what happens in this situation, possibly because it has not occurred until now. The design assumes that transactions on different cache lines can progress independently.
In most implementations, this assumption typically holds in order to maximize the memory bandwidth utilization and minimize request latency, and also to simplify reasoning about deadlock freedom of the coherence protocol itself. However, it may be the case that the cache stripes transactions across a limited number of independent units and which might deadlock if both A and B were mapped to the same unit. Were this to be the case, it could be avoided by careful placement of A and B in the physical address space. This is actually the case in our FPGA device implementation, described in Section 5.
Implementation in Enzian
We implemented the two device-homed protocols over Enzian and evaluate them in Section 6. The implementations use SystemVerilog and SpinalHDL (contributors, 2024) and exploit an existing “Directory Controller” on Enzian’s FPGA (Ramdas, 2023) which manages the coherent links to the CPU and handles transient states in the ECI protocol.
ECI provides all the messages we require to implement our messaging protocols. In particular, it includes the ability for the device to return a line to the CPU cache in Exclusive state, even though the CPU requested it in Shared. This optimization dramatically reduces latency, as shown in Section 6.1.
In our case, no flow-control issues arise since ECI uses separate virtual channels for request and reply messages, removing the chance of deadlock. However, other complications arise: ThunderX-1 generates Global Synchronization (GSync) messages which function as global barriers. A GSync blocks any subsequent transactions until all those started before the GSync have retired. Fortunately, any invalidations sent are not blocked by core, allowing the protocol to progress regardless.
The ThunderX-1 divides its last-level cache functionality into a set of units called TADs, each of which can handle up to 16 simultaneous cache transactions. We ensure that consecutive cache lines are mapped to different TADs to ensure that the operations can proceed independently.
@@ 6. Evaluation
We compare PIO over a coherent interconnect with two other reference points: PIO directly over PCIe– essentially, the CPU copying data to and from a region of PCIe address space – and descriptor-based DMA over PCIe. All use the same Enzian hardware platform. Coherent PIO uses the Enzian native ECI, while the two PCIe techniques connect the CPU and FPGA PCIe interfaces back-to-back with a loopback cable as described in Section 3, effectively turning the Enzian FPGA into a conventional PCIe FPGA card. DMA experiments all use the Xilinx XDMA hardware IP on the FPGA.
@@
6.1. Invoking functions on an accelerator
We first measure latency when the CPU invokes a function on the device, passing arguments, and obtaining a result. We use the protocol shown in 5(a), whereby the cache lines are homed on the FPGA.
Figure 6. Invocation latency over ECI vs. FastForward
Figure 6 shows the distribution of total latency for the this operation. We see a median latency of around 1600ns for the unoptimized protocol, wherein the result line is returned in Shared state as it was requested by the CPU. When we introduce the optimization of returning the line in Exclusive, the median latency drops to around 900ns.
For comparison, we also show an implementation of the FastForward protocol (Giacomoni et al., 2008) exchanging cache lines between CPU sockets in a dual-socket ThunderX-1-based Gigabyte R150-T61 server with similar CPU and DDR specification to Enzian; this achieves a median latency of about 1150ns.
The optimized protocol is performing 2 round-trip message exchanges over ECI, which has a one-way latency at the link layer of about 150ns. The rest of the overhead (300ns) is mostly incurred in the protocol processing in the directory controller. The FPGA is clocked at about 300MHz.
A significant penalty is incurred when the results of the call are returned in Shared, resulting in further round trips over ECI to (unnecessarily) invalidate the line at the FPGA before sending the next message. This suggests that future coherent interconnects would benefit greatly from allowing the home node to return a line in Exclusive.
Overall, the optimized protocol comfortably outperforms FastForward, despite the latter exploiting highly optimized hardware implementations at both ends (clocked at 2.1GHz) and aggressively polling, whereas our optimized protocol requires no spinning on the part of the CPU or device.
6.2. Invocation over ECI vs. PCIe PIO and DMA
@@ 6.1. Invoking functions on an accelerator
We now compare the latency achieved in the previous section with both PIO over PCIe and the Xilinx XDMA hardware using descriptors. For DMA numbers we use the Xilinx XDMA benchmark (Xilinx DMA Benchmark, 2024), which essentially performs the same operation using descriptor-based DMA. In both these cases, the FPGA is configured to map the request and response into a write to, and a read from, onboard Block RAM.
@@
As in Section 6.1, we measure latency to send a request from the CPU to the FPGA and read a response. However, this time we vary the payload size (of both request and response) and report median latency figures.
@@
The results are shown in Figure 7. Up to payloads of 8KiB, the latency of XMDA-based communication is dominated by descriptor setup and manipulation, and remains largely independent of payload size. Faster PCIe interconnects (this is a PCIe Gen3 x8 link) would amplify this effect.
Latency over ECI is dramatically lower for smaller payloads. Unsurprisingly, it is constant up to 128 bytes (the size of a cache line on Enzian) and then increases linearly, since the protocol simply replicates each cache line exchange until all the data has been transferred. It is only at 8KiB that ECI invocation latency exceeds that of XMDA. Arguably the transfer of multiple cache lines over ECI might be improved by clever pipelining and locating the lines in physical memory so as to exploit parallelism in the directory controllers, but we do not explore that in this paper.
Relative to ECI, PIO over PCIe performs poorly except for payloads under 16 bytes in size. This is due to limitations in PCIe that we have discussed in Section 4, notably the limited read bus width.
We conclude that up to the page size (4KiB) or the typical size of jumbo frames, PIO remains significantly faster than the DMA option commonly used for these transaction sizes. Thus even using the XDMA implementation for the experiments we still have the room for comparison with optimized DMA implementations.
This experiment is closed-loop and single-threaded, and so achieved throughput is proportional to the round-trip latency. Fully saturating the raw link bandwidth (in all cases) would require many parallel invocations at once, which would go against our goal of supporting applications where latency is paramount.
@@ 6.3. Network interface adapter
We implement 3 variations of a 100 Gb/s NIC in Enzian. The first (8(a)) is a conventional approach connecting one of Enzian’s 100 Gb/s Ethernet MACs directly to the XDMA engines, resulting in a NIC programmed using descriptor rings. The second design (8(b)) buffers packets in onboard SRAM on the FPGA and allows the CPU to read and write packet and control data using PCIe reads and writes. The final approach (8(c)) uses a custom module on the FPGA to bridge between the ECI directory controller and the Ethernet MAC, using a slight variant of the protocol in 5(b) to deliver network packets to the CPU in multiple cache lines. In all cases we clock the core NIC logic at 250 MHz.
Figure 9. NIC implementations: receive latency
Figure 10. NIC implementations: transmit latency
@@
For experiments we configure the Ethernet MAC in near-end PCS/PMA loopback mode for a reliable measurement of packet delivery inside the server without network delays. We define the receive latency as time passed from the last beat of the packet appearing on the ingress streaming interface of the Ethernet MAC, to the CPU fully receiving the packet content in its registers. The transmit latency is defined similarly as time passed from the CPU having the packet ready in registers, to the last beat of the packet appearing on the egress streaming interface on the MAC. We do not include the time it takes the packet to go through the Ethernet PCS/PMA loopback, since the MAC is fixed and the same in all implementations presented in this experiment.
@@
To minimize disruptions from the Linux scheduler, we follow the common practice to free one CPU core from kernel tasks and IRQ processing (Akkan et al., 2012), with the isolcpus function together with core pinning using taskset. In this setup, the only disturbance from Linux would be the 250 Hz timer interrupt present on all cores. While a custom, completely tickless kernel would remove this disturbance completely, it is not commonly deployed in production environments due to performance complications. We use such a tickless kernel to measure tail latency below, but other experiments use a stock generic kernel from Ubuntu.
We show figures for XDMA in both polled and interrupt-driven modes; in practice the difference in latency is small.
@@
Figure 9 shows receive latency for different packet sizes for each of the NIC implementations. As in the invocation experiment, DMA latency is fairly stable across all packet sizes, suggesting that the latency is dominated by various DMA overheads, such as descriptor setup and manipulation.
Note that system calls here actually result in little overhead (only a few $\mu$s); the overhead is dominated by cache misses manipulating DMA data structures in main memory.
@@
In contrast, PIO over both ECI and PCIe offer much lower latency for small packet sizes up to 1024 bytes DMA, highlighting high efficiency gains in PIO compared to DMA. However, the latency of PCIe PIO quickly degrades for larger packets due to the non-posted read problem discussed in Section 4, reaching over 350 $\mu$s for 8 KiB packets. In contrast, ECI offers dramatically better latency even at 8 KiB packets.
@@
Figure 10 shows transmit latency. In this case, PCIe PIO competes well with ECI across all packet sizes due to combining posted writes as we discussed in Section 4. PIO over ECI shows a slightly higher latency for small packets, due to it needing 2 round-trips over ECI for each 128 bytes, versus a single PCIe round trip. Both PIO solutions perform significantly better than DMA across all packet sizes.
@@
Table 1 shows the tail latency impact for selected packet sizes for each of the NIC implementations. We select three representative packet sizes to evaluate tail latency: 64 for the smallest possible packet, 1536 for the normal Ethernet MTU, and 9600 for a common jumbo frame MTU. We use the tickless kernel setup mentioned earlier to avoid the 250 Hz Linux timer, in order to reveal the actual tail latency impact from our NIC.
@@
We see that PIO over PCIe results in low tail latency compared to the descriptor-based approach, while ECI completely eliminates tail latency in this scenario.
@@
These results show a clear advantage of using PIO over DMA, even for larger packets, and especially when using a coherent interconnect . The poor receive performance of PCIe-based PIO for receiving packets is due to the unposted-read limitation of PCIe, and disappears for transmit.
6.4. Accelerating Timely Dataflow
In this example we apply the techniques and results outlined so far to a practical workload: Timely Dataflow (Murray et al., 2013), a real-time stream-processing system. We compare the baseline performance to a partially FPGA-offloaded implementation under three CPU–FPGA communication scenarios: DMA over PCIe, PIO over PCIe, and PIO over ECI.

### v3

@@ 4. PIO over a coherent interconnect
    The CPU cache receives $A$ and the CPU unblocks, immediately loading the first word of the response into a register.
This would lead to $A$ in Shared in both nodes, since software requested $A$ using a load. However, as an optimization, the device can return the line in Exclusive instead, invalidating its own copy, so the protocol returns to a quiescent state with the roles of the lines reversed.
This protocol either eliminates or minimizes the drawbacks of a software-only protocol: the ability to stall the CPU eliminates both the need to poll, and the race condition. By coupling and overlapping two coherence transactions we are able to transfer data in both directions with the same two round-trips. All data is transferred atomically so there is no risk of half a payload being shipped. This is the best achievable result with an unmodified coherence protocol, since no cache transaction both sends and receives a payload.
Figure 6. Invocation latency over ECI vs. FastForward
Figure 6 shows the distribution of total latency for this operation implemented on Enzian. Median latency is around 1600ns when the result is returned in Shared state as requested by the CPU (“ECI unopt”). When we return the line in Exclusive, this drops to around 900ns (“ECI”).
For comparison, we also show an implementation of the FastForward protocol (Giacomoni et al., 2008) exchanging cache lines between CPU sockets in a dual-socket ThunderX-1-based Gigabyte R150-T61 server with similar CPU and DDR specification to Enzian; this achieves a median latency of about 1750ns.
The optimized protocol is performing 2 round-trip message exchanges over ECI, which has a one-way latency at the link layer of about 150ns. The rest of the overhead (300ns) is mostly incurred in the protocol processing in the directory controller, and would be lower in an ASIC implementation. The FPGA is clocked at about 300MHz.
Returning a line in Exclusive
The ability for the device to unilaterally return a line in Exclusive delivers an important performance benefit. The ECI protocol, and the ThunderX-1 last-level cache, both support this. Interestingly, it is also proposed in the (as yet unimplemented) CXL.mem 3.0 standard (CXL Consortium, 2024), along with the “back invalidate” operation providing symmetric cache coherence. We suggest this is a useful feature for future protocols.
Handling larger messages
The protocols above handle one line at a time (128B on Enzian). They can be used repeatedly to transfer more data than this, but this is inefficient. Instead, we extend the protocols described so far using parallelism and pipelining to give much better throughput. This leads to two further protocols, which are the ones we evaluate in Section 5.
For the NIC example of Section 5.2, with predominantly unidirectional communication, we dedicate one pair of cache lines as control (as in 5(a) and 5(b)) and add overflow cache lines to carry data. When the protocol invalidates the control lines, it also invalidates these overflow cache lines in parallel, exploiting the full bandwidth of the coherent link. We use this technique to transfer packets larger than one cache line, up to an Ethernet jumbo frame (9kB).
We can do even better when the application has a symmetrical communication pattern, i.e. the CPU sends and receives similar amounts of data in one invocation. This is the case for the device invocation and Timely Dataflow experiments we show in Section 5.1 and Section 5.3. In this case, the better option is to prefetch multiple instances of the protocol shown in 5(c).
To transfer $n$ cache lines at a time, we maintain two groups of $n$ lines. The CPU first writes all arguments for the device to the lines in group 0, and then prefetches all cache lines in group 1. The prefetches trigger the device logic to invalidate all cache lines in group 0, fetching the arguments. The prefetches are issued in parallel by the CPU and again provide enough concurrency to saturate the link.
Handling timeouts
The protocols block a core’s request for a cache line until a device operation has completed. This timeout is typically quite long (often hundreds of milliseconds) but if it happens the CPU’s cache is likely to cause an exception, or other fault like a machine check.
@@
Avoiding deadlocks
It is also important to prevent deadlocks due to the device stalling a request until it can issue, and get a reply back for, another request to a different cache line. Existing standards for inter-operable cache coherent interconnects are silent on what happens in this situation, possibly because it has not occurred until now. The design assumes that transactions on different cache lines can progress independently.
In some implementations, this assumption holds in order to maximize the memory bandwidth utilization and minimize request latency, and also to simplify reasoning about deadlock freedom of the coherence protocol itself. However, it may be the case that the cache stripes transactions across a limited number of independent units which might deadlock if both $A$ and $B$ were mapped to the same unit. Were this to be the case, it could be avoided by careful placement of $A$ and $B$ in the physical address space.
This is actually the case in the Enzian FPGA directory controller implementation we use (Ramdas, 2023), which follows similar provisioning choices in the CPU. The ThunderX-1 divides its last-level cache functionality into a set of units called TADs, each of which can handle up to 16 simultaneous cache transactions. We ensure that consecutive cache lines are mapped to different TADs to ensure that the operations can proceed independently.
Another possible source of deadlocks is that the CPU and L2 cache might issue requests out of order, especially in the case of using prefetches. This is mainly relevant for handling requests larger than one cache line, as we discussed above. The device implementation needs to be aware of possible reorderings and should not rely on the request ordering to perform state transition to avoid deadlocks. In our implementation, we always buffer possible in-flight cache lines on the FPGA and advance state machines based on number of requests we see, without relying on the actual order of them.
Correctness and interoperability
An obvious concern given the complexity of underlying cache coherence protocols is the correctness of the implementation. An error in the device might cause the state machine in the CPU’s cache to enter a state from which it cannot recover.
Worse, while our implementation relies on an existing, and well-tested, directory controller implementation on the FPGA, we are using the ECI protocol in unconventional ways unforeseen by the designers of the CPU and its cache.
In practice, after extensive testing we have not seen protocol-related problems even under heavy load. Moreover, we are reassured by recent work on exhaustive semi-open-state testing of coherence protocols (Schult et al., 2024) which can formally define the state space of a subset of a coherence protocol (which can nevertheless include the kinds of behaviors we describe here), and then exhaustively test a hardware implementation against that formal model.
Enzian-specific implementation issues
Our implementations will all be made available as open-source and use a combination of SystemVerilog and SpinalHDL (contributors, 2024), and use a version of the existing Enzian FPGA directory controller (Ramdas, 2023).
@@
Generality
If we had only simulated the protocols in Figure 5 rather than implement them in a hardware platform, it seems unlikely that we could have become aware of the numerous subtleties and challenges that real coherent interconnects present. However, this raises generality concerns, since our implementation is tied to a single platform.
In principle, all the protocols we have described here could be implemented over any coherent interconnect which uses a symmetric, MOESI-like, distributed directory-based model of coherence, and where the interconnect hardware exposed major state transitions to the device-specific logic which served as the end-point for messaging.
Several current and future standards satisfy these conditions, for example CCIX (CCIX Consortium and others, 2024), TileLink (Berkeley Architecture Research, 2022), and CXL.mem 3.0 (CXL Consortium, 2024), although notably neither CXL 2.0 nor CXL.cache 3.0. We are also seeing something of a boom in custom interconnects from a range of vendors. By making the requirements for this fine-grained, low-latency communication clear, future designers can take this into account when designing new interconnects and parts.
@@ 5. Evaluation
We compare coherent PIO with our two other reference points: optimized PIO over PCIe described in the previous section, and descriptor-based DMA over PCIe. All use the same Enzian platform. Coherent PIO uses the Enzian native ECI implementation, while the PCIe approaches use a loopback cable between PCIe interfaces on the CPU and FPGA, effectively turning the Enzian FPGA into a conventional PCIe FPGA card. DMA experiments use the Xilinx XDMA IP.
@@
With the exception of the NIC case, the results show very little variance. All graphs show 95% confidence intervals, but these are typically not visible.
5.1. Accelerator invocation
@@
We first explore the scenario where the CPU invokes a function on the device configured as an accelerator, using the three communication approaches to send argument data and receive result data. In this case, the FPGA is configured to map the invocation into a write to, followed by a read from, onboard Block RAM.
For DMA numbers we use the Xilinx XDMA benchmark (Xilinx DMA Benchmark, 2024), which performs the operation using descriptor-based DMA. We implement coherent PIO with the protocol shown in 5(c) to send a request and receive a response in two round-trips.
Latency comparison
@@
We first compare the latency for the CPU to complete a function invocation on the device, measured as the time taken for the CPU to send a request from the CPU to the FPGA and read a response. We vary the payload size (of both request and response) and report median latency figures.
@@
The results are shown in Figure 7. Up to payloads of 8KiB, the latency of DMA-based communication is dominated by descriptor setup and manipulation, and remains largely independent of payload size (upper scale). Faster PCIe interconnects or CXL would only amplify this effect.
Latency over ECI is dramatically lower across the board. It is constant up to 256 bytes (two Enzian cache lines) and then increases slowly, but linearly. This is due to the pipelined nature of the ECI PIO protocol: before we saturate the links, the incremental latency induced by each extra line is small.
Relative to ECI, PIO over PCIe performs poorly for any payload more than 16 bytes, due to the inherently-limited PIO read bandwidth discussed in Section 3, but competes well with DMA up to 2KiB payload.
We conclude that, for almost all transfers up to and beyond 8 KiB, coherent PIO is significantly lower latency than both the DMA option commonly used for these transaction sizes, and PIO over PCIe. More PCIe bandwidth (e.g. PCIe Gen5) will not change the PIO results significantly without support for more efficient MMIO reads, since PCIe round-trip latency remains the same. Neither would it improve DMA latency, which is dominated by descriptor overhead. In contrast, the lower interconnect latency available with newer CXL versions would improve things, but would also deliver the same benefit to the coherent PIO case.
Figure 8. Invocation throughput for different payload sizes
Throughput comparison
We also show throughput figures, omitting PIO over PCIe since performance (as seen above) is not practical for larger transfers. We measure how many back-to-back invocations one CPU core can make for 1 ms, using different transaction sizes. We calculate throughput by the total data volume transferred divided by the time elapsed and report the median value in Figure 8.
ECI PIO throughput increases with transaction size up to 32KiB payloads to a peak of 2.19 GiB/s and then dropping slightly, due to thrashing in the 32KiB L1 data cache. It performs better than PCIe DMA over all transaction sizes presented, and by a comfortable margin. At 64 KiB, we are still far from saturating throughput for PCIe DMA; it is well known that very large transaction sizes are needed to achieve maximum throughput.
In conclusion, ECI PIO shows a clear throughput advantage over PCIe DMA at multiple page sizes (up to 64KB) for device invocations, to the extent that the determining factor is less the throughput achieved and more the number of CPU cycles or cache size available.
As with the latency, this is unlikely to change with faster PCIe interconnects, and CXL would likely result in improvements across the board, including for the coherent PIO.
The results also show that there is an optimal native transaction size for coherent PIO, in this case the L1 cache size: larger transfers should be broken down into smaller transactions of optimal size to achieve maximum throughput.
@@ 5.2. Network interface adapter
We implement 3 variations of a 100 Gb/s NIC in Enzian. The first (9(a)) is a conventional approach connecting one of Enzian’s 100 Gb/s Ethernet MACs directly to the XDMA engines, resulting in a NIC programmed using descriptor rings. The second design (9(b)) buffers packets in onboard SRAM on the FPGA and allows the CPU to read and write packet and control data using PCIe reads and writes. The final approach (9(c)) uses a module on the FPGA to bridge between the ECI directory controller and the MAC. A variant of the protocol in 5(b) delivers network packets to the CPU in multiple cache lines, and 5(a) is used to send a packet. The NIC logic is clocked at 250 MHz.
Figure 10. Latency of NIC implementations
@@
For experiments we configure the Ethernet MAC in near-end PCS/PMA loopback mode to reliably measure packet delivery inside the server without network delays. We define receive latency as time from the last beat of the packet appearing on the ingress streaming interface of the MAC, to the CPU fully receiving the packet contents in its registers. Transmit latency is similarly defined as time from the CPU having the packet ready in registers, to the last beat of the packet appearing on the egress streaming interface on the MAC. We do not include the time for the packet to go through the Ethernet PCS/PMA loopback, since the MAC is fixed and the same in all implementations presented.
@@
To minimize disruptions from the Linux scheduler, we follow the common practice to free one CPU core from kernel tasks and IRQ processing (Akkan et al., 2012), with the isolcpus function together with core pinning using taskset. In this setup, the only disturbance from Linux would be the 250 Hz timer interrupt present on all cores. While a custom, completely tickless kernel would remove this disturbance completely, it is not commonly deployed in production environments due to performance complications. We use such a tickless kernel to measure tail latency below, but other experiments use a stock generic kernel from Ubuntu. We show figures for XDMA in both polled and interrupt-driven modes; in practice the difference in latency is small.
@@
Figure 10 shows receive latency for different packet sizes for each NIC implementation. As before, DMA latency is stable across all packet sizes, suggesting that it is dominated by various DMA overheads, such as descriptor setup and manipulation. Note that system calls here actually result in little overhead (only a few $\mu$s); the overhead is dominated by cache misses manipulating DMA data structures in RAM.
@@
In contrast, PIO over both ECI and PCIe offer much lower latency for small packet sizes up to 1024 bytes DMA, highlighting high efficiency gains in PIO compared to DMA. However, the latency of PCIe PIO quickly degrades for larger packets due to the non-posted read problem discussed in Section 3, reaching over 350 $\mu$s at 8 KiB packets. In contrast, ECI offers dramatically better latency even at this size.
@@
Figure 10 also shows transmit latency. PCIe PIO competes well with ECI across all packet sizes due to combining posted writes. PIO over ECI has slightly higher latency for small packets, since it needs 2 round-trips over ECI for each 128 bytes, versus a single PCIe round trip. Both PIO solutions perform significantly better than DMA across all packet sizes.
@@
Table 1 shows the tail latency impact for selected packet sizes for each NIC implementation. We select three representative packet sizes to evaluate tail latency: 64 for the smallest possible packet, 1536 for the normal Ethernet MTU, and 9600 for a common jumbo frame MTU. We use the tickless kernel setup mentioned earlier to avoid the 250 Hz Linux timer, in order to reveal the actual tail latency impact from our NIC.
@@
PIO over PCIe results in low tail latency compared to the descriptor-based approach, while ECI completely eliminates tail latency in this scenario.
@@
These results show clear benefits of using PIO vs. DMA, even for large packets, and especially when using a coherent interconnect. The poor receive performance of PCIe-based PIO for receiving packets is due to the unposted-read limitation of PCIe, and disappears for transmit. This advantage should translate to future CXL3.0 and PCIe Gen6 NICs.
5.3. Offloading Timely Dataflow